# Automated Risk Management Mechanisms in DeFi Lending Protocols: A Crosschain Comparative Analysis of Aave and Compound

### v1

@@ Abstract
Blockchain-based decentralised lending is a rapidly growing and evolving alternative to traditional lending, but it poses new risks. To mitigate these risks, lending protocols have integrated automated risk management tools into their smart contracts. However, the effectiveness of the latest risk management features introduced in the most recent versions of these lending protocols is understudied. To close this gap, we use a panel regression fixed effects model to empirically analyse the cross-version (v2 and v3) and cross-chain (L1 and L2) performance of the liquidation mechanisms of the two most popular lending protocols, Aave and Compound, during the period Jan 2021 to Dec 2024. Our analysis reveals that liquidation events in v3 of both protocols lead to an increase in total value locked and total revenue, with stronger impact on the L2 blockchain compared to L1. In contrast, liquidations in v2 have an insignificant impact, which indicates that the most recent v3 protocols have better risk management than the earlier v2 protocols. We also show that L1 blockchains are the preferred choice among large investors for their robust liquidity and ecosystem depth, while L2 blockchains are more popular among retail investors for their lower fees and faster execution.
Index Terms:
Blockchain, DeFi, Lending, Risk Management
@@ I Introduction
The remainder of this study is arranged as follows: Section II presents an introduction to the DeFi lending market and discusses the relevant literature in this field and how our research contributes to it. In Section III, we present our data collection process. In Section IV, we discuss the empirical methodology used in our study. Section V interprets the relevant empirical findings and discusses the corresponding results. Finally, Section VI concludes.
@@ II-A Evolution of DeFi lending Market, Protocols, and Advanced Risk Management
Table I summarizes the functionality of Aave and Compound, two of the oldest and most successful DeFi lending platforms. Aave has the highest TVL in the DeFi lending space and is a decentralised open-source platform operating on 11 different blockchains. Compound finance is also among the top five lending protocols and operates on nine different blockchains. Notably, initial versions (i.e., v1, v2) of these protocols only implemented basic risk controls, whereas upgraded versions (i.e., v3) introduced more sophisticated tools that are explicitly designed to mitigate systemic vulnerabilities. These upgraded versions are launched on both L1 and L2 blockchains: L1 blockchains (e.g., Ethereum) denote the base networks that establish the primary consensus processes to achieve decentralization and guarantee primary on-chain activities; while L2 blockchains (e.g., Arbitrum, Optimism) inherit security and decentralization from L1 and increase transaction throughput, offering low transaction fees and faster transactions [20].
@@
Risk management in lending protocols relies on automated mechanisms like over-collateralisation ratios, governance frameworks, and liquidation triggers to protect participants and maintain stability. However, these measures can be insufficient during periods of high volatility, market congestion, liquidation spirals, and systemic risk. High transaction costs can also reduce their effectiveness in maintaining stability [16, 17]. The latest v3 versions of the protocols address these shortcomings by incorporating advanced risk management features, as shown in Table I; including enhanced liquidation incentives and dynamic collateral requirements, which allow users to adjust their borrowing parameters to match their risk profile, thus mitigating liquidation cascades and improving resilience. Furthermore, the integration of L2 blockchain with v3 protocols has reduced transaction cost and increased throughput, thus facilitating faster and more efficient collateral modifications and immediate liquidation triggers, which mitigates the prospect of delayed or unsuccessful liquidations. As more than 70% of liquidations occur in the absence of any price decline of crypto assets [22], the latest protocols also observe market sentiment and patterns for efficient risk management. These innovations collectively facilitate a more resilient and efficient risk management system (particularly when compared with v2 on L1 blockchain).
@@ III Data Collection
DRM=Loan-To-ValueLiquidation Threshold(1+Liquidation $\text{DRM}=\frac{\text{Loan-To-Value}}{\text{Liquidation Threshold}}\;\left(1+\text{Liquidation Penalty}\right)$ (1)
@@ IV Methodology
Regression analysis is the most common method to determine the relationship between dependent and independent variables in DeFi markets. Studies that have used regression analysis to assess various aspects and risks in DeFi lending markets include [2, 3, 12, 16, 17, 19, 22, 25]. In this study, we employ a panel regression model with fixed effects to quantify how liquidations—driven by various risk factors—affect the performance of DeFi lending protocols. The regression model is illustrated in Equation (IV).
@@
The dependent variable, $\text{Perf}_{i,t}$, measures the performance of protocol $i$ at time $t$, using either TVL or TR data after logarithmic transformation. Our main explanatory variable is the logged liquidation volume, $\mathrm{Liq}_{i,t}$. We include several control variables to account for potential confounding factors. The variable $\text{DRM}_{i,t}$ captures the level of default risk associated with protocol $i$ at time $t$. Market volatility is represented by $\text{VolETH}_{i,t}$ for the collateral asset ETH, $\text{VolGasPri}_{i,t}$ for the volatility of gas prices, and $\text{VolUSDCbor}_{i,t}$ for the volatility of USDC borrow rate. The volatilities are 7-day rolling standard deviation of the log-returns. Protocol-level activity is measured by $\text{Wd}{i,t}$, which is the logged total withdrawal volume from protocol $i$ at time $t$. To control for broader market sentiment, we incorporate $\text{FGI}_{i,t}$ as a proxy for investor sentiment in the cryptocurrency market and $\text{VIX}_{i,t}$ for sentiment in the traditional financial market. The sentiment variables are used after logarithmic transformation.
[…]
@@ V Results and Findings
V-A Cross-version and cross-chain analysis of liquidations
It should be noted that the coefficient $\mathrm{Liq\mathord{:}Lay}$ is significantly positive at $0.05$, indicating that liquidation has a heterogeneous impact on performance across different layers under the same version, specifically in v3. This suggests that a $1\%$ increase in liquidation on Arbitrum is associated with a $0.05\%$ greater increase in TVL compared to Ethereum, highlighting a differentiated response to liquidation activity between the two layers. This is a new finding that has not been reported before.
@@ V-A Cross-version and cross-chain analysis of liquidations
Liq $-0.01$∗∗ $0.00$ $0.00$ $0.00$
Liq:Lay $0.05$∗∗ $0.09$∗∗
Liq:Ver $0.03$∗∗ $0.02$  $0.05$∗∗ $0.06$∗∗
Liq:Platf $-0.01$  $0.00$
DRM $0.01$∗∗ $0.01$∗∗ $0.01$∗∗ $0.01$∗∗
VolETH $-1.90$∗∗ $-2.23$∗ $-1.94$∗∗ $-2.25$∗
VolGasPri $-0.15$∗∗ $-0.19$∗∗ $-0.15$∗∗ $-0.19$∗∗
VolUSDCbor $0.02$∗∗ $0.05$∗∗ $0.02$∗∗ $0.05$∗∗
Wd $0.40$∗∗ $0.49$∗∗ $0.40$∗∗ $0.49$∗∗
FGI $0.08$∗∗ $0.46$∗∗ $0.09$∗ $0.48$∗∗
VIX $-0.27$∗∗ $-0.76$∗∗ $-0.29$∗∗ $-0.79$∗∗
@@ VI Conclusion
We have investigated the effectiveness of automated risk management strategies during liquidation events and their impact on the performance and stability of decentralised lending protocols. For this purpose, we selected the top lending protocols, Aave and Compound, including their old (v2) and upgraded (v3) versions deployed on L1 and L2 blockchains, to assess the effectiveness of prevalent risk management mechanisms. The results suggest that liquidation events in v3 protocols lead to increased TVL and total revenue, with a more pronounced effect on the Arbitrum (L2) blockchain compared to the Ethereum (L1) blockchain. In contrast, for the v2 protocols, liquidations have an insignificant impact on TVL and total revenue. A combination of protocol-specific and market-driven factors is likely to influence the behavioral divergence between v2 and v3. Our research has important implications for risk managers, investors, and policy makers.
This research could be further extended by investigating the wallet-level behavior of lenders, borrowers, and liquidators during and after liquidation events to assess strategic user actions. Furthermore, an event study could be used to explore the role of specific stablecoins and volatile assets in causing systemic risk or spillover effects, after cascading liquidations. Another avenue of investigation is to explore the total value redeemable (TVR) as a performance measure alternative to TVL, as it has been proposed as a more accurate quantifier of the value stored within a protocol [11].
@@ Acknowledgements
E. Iftikhar’s research is supported by PNRR (National Recovery and Resilience Plan) scholarship ex D.M. 351/2022 under National Call for PhD in Blockchain and Distributed Ledger Technology offered by University of Camerino (Administrative University) and Universita Cattolica del Sacro Cuore, Milan (Host University). This work was completed at the University of Bristol while E. Iftikhar was a visiting PhD scholar, under the supervision of W. Wei and J. Cartlidge.
@@ Appendix A Multicollinearity Test
Variable VIF
Liq 1.25
DRM 1.13
VolETH 1.28
VolGasPri 1.61
VolUSDCbor 1.61
FGI 2.11
VIX 1.84
Wd 1.36
@@ Appendix B Comprehensive Model
Here we have used all variable mentioned in our regression model (see Section IV) to account for cross-version, cross-chain and cross-protocol affection on protocol performance. Our findings in Table B.2 are consistent with the conclusions we made in Section V.
The results show that there is significant heterogeneity between layers and versions, which is consistent with the results of Models 1–4.
@@
Liq:Lay $0.05$∗∗ $0.09$∗∗
Liq:Ver $0.03$∗∗ $0.02$
Liq:Platf $-0.01$  $0.00$
DRM $0.01$∗∗ $0.01$∗∗
VolETH $-1.93$∗ $-2.23$∗
VolGasPri $-0.15$∗∗ $-0.19$∗∗
VolUSDCbor $0.02$∗∗ $0.05$∗∗
Wd $0.40$∗∗ $0.49$∗∗
FGI $0.08$∗ $0.46$∗∗
VIX $-0.27$∗∗ $-0.76$∗∗

### v2

@@ Abstract
Blockchain-based decentralised lending is a rapidly growing and evolving alternative to traditional lending, but it poses new risks. To mitigate these risks, lending protocols have integrated automated risk management tools into their smart contracts. However, the effectiveness of the latest risk management features introduced in the most recent versions of these lending protocols is understudied. To close this gap, we use a panel regression fixed effects model to empirically analyse the cross-version (v2 and v3) and cross-chain (L1 and L2) effectiveness of liquidation mechanisms, measured through TVL and total revenue as proxies for performance of the two most popular lending protocols, Aave and Compound, during the period Jan 2021 to Dec 2024. Our analysis reveals that liquidation events in v3 of both protocols lead to an increase in total value locked and total revenue, with stronger impact on the L2 blockchain compared to L1. In contrast, liquidations in v2 have an insignificant impact, which indicates that the most recent v3 protocols have better risk management than the earlier v2 protocols. We also show that L1 blockchains are the preferred choice among large investors for their robust liquidity and ecosystem depth, while L2 blockchains are more popular among retail investors for their lower fees and faster execution.
@@ I Introduction
The remainder of this study is arranged as follows: Section II presents an introduction to the DeFi lending market and discusses the relevant literature in this field and how our research contributes to it. In Section III, we present our data collection process. In Section IV, we discuss the empirical methodology used in our study. Section V interprets the relevant empirical findings and discusses the corresponding results. Finally, Section VI concludes. All code and data are publicly available. [^1] https://github.com/ErumIftikhar/DefiLending_Data_Code.
@@ II-A Evolution of DeFi lending Market, Protocols, and Advanced Risk Management
Table I summarizes the functionality of Aave and Compound, two of the oldest and most successful DeFi lending platforms. Aave has the highest TVL in the DeFi lending space and is a decentralised open-source platform operating on 11 different blockchains. Compound finance is also among the top five lending protocols and operates on nine different blockchains. Notably, initial versions (i.e., v1, v2) of these protocols only implemented basic risk controls, such as collateral factors and static risk parameters (liquidation threshold, liquidation penalties, borrow/lending rates), whereas upgraded versions (i.e., v3) introduced more sophisticated tools that are explicitly designed to mitigate systemic vulnerabilities. These tools differ from traditional risk management practices (ring-fencing, credit exposure limits, and portfolio rebalancing) by being automated, transparent, and enforced via smart contracts. The dynamic risk parameters and targeted mechanisms provide more granular control over asset-specific risk and reduce the likelihood of systemic contagion. These upgraded versions are launched on both L1 and L2 blockchains: L1 blockchains (e.g., Ethereum) denote the base networks that establish the primary consensus processes to achieve decentralization and guarantee primary on-chain activities; while L2 blockchains (e.g., Arbitrum, Optimism) inherit security and decentralization from L1 and increase transaction throughput, offering low transaction fees and faster transactions [21].
Risk management in lending protocols relies on automated mechanisms like over-collateralisation ratios, governance frameworks, and liquidation triggers to protect participants and maintain stability. However, these measures can be insufficient during periods of high volatility, market congestion, liquidation spirals, and systemic risk. High transaction costs can also reduce their effectiveness in maintaining stability [16, 17]. The latest v3 versions of the protocols address these shortcomings by incorporating advanced risk management features, as shown in Table I; including enhanced liquidation incentives and dynamic collateral requirements, which allow users to adjust their borrowing parameters to match their risk profile, thus mitigating liquidation cascades and improving resilience. Isolation mode prevents contagion from volatile assets; efficiency mode, optimizes capital utilization for correlated assets; and caps on borrow and supply asset limits, overexposure to risky tokens. Furthermore, the integration of L2 blockchain with v3 protocols has reduced transaction cost and increased throughput, thus facilitating faster and more efficient collateral modifications and immediate liquidation triggers, which mitigates the prospect of delayed or unsuccessful liquidations. As more than 70% of liquidations occur in the absence of any price decline of crypto assets [22], the latest protocols also observe market sentiment and patterns for efficient risk management. These innovations collectively facilitate a more resilient and efficient risk management system (particularly when compared with v2 on L1 blockchain). Our study focuses on investigating those risk parameters that exert a direct impact on the liquidation risk mechanism, such as to loan-to-value ratio, liquidation penalty, and liquidation threshold.
@@ III Data Collection
$\text{DRM}=\frac{\text{Loan-To-Value}}{\text{Liq. Threshold}}\;\left(1+\text{Liq. Penalty}\right)$ (1)
@@ IV Methodology
Regression analysis is the most common method to determine the relationship between dependent and independent variables in DeFi markets. Studies that have used regression analysis to assess various aspects and risks in DeFi lending markets include [2, 3, 12, 16, 17, 19, 22, 25]. In this study, we employ a panel regression model with fixed effects to quantify how liquidations—driven by various risk factors—affect the performance (proxied by TVL and TR) of DeFi lending protocols. As our dataset spans multiple protocols (Aave and Compound, across v2/v3 and L1/L2 deployments) over time, panel regression analysis is well-suited because it captures both the cross-sectional variation between protocols and the time-series variation within each protocol. Panel regression with fixed effects controls for unobserved, time-invariant heterogeneity and improves estimation efficiency by leveraging repeated observations over time [26]. The regression model is illustrated in Equation (IV).
@@
The dependent variable, $\text{Perf}_{i,t}$, measures the performance of protocol $i$ at time $t$, using either TVL or TR data after logarithmic transformation. Our main explanatory variable is the logged liquidation volume, $\mathrm{Liq}_{i,t}$. We include several control variables to account for potential confounding factors. The variable $\text{DRM}_{i,t}$ captures the level of default risk associated with protocol $i$ at time $t$. Market volatility is represented by $\text{VolETH}_{i,t}$ for the collateral asset ETH, $\text{VolGasPri}_{i,t}$ for the volatility of gas prices, and $\text{VolUSDCbor}_{i,t}$ for the volatility of USDC borrow rate. The volatilities are 7-day rolling standard deviation of the log-returns. Protocol-level activity is measured by $\text{Wd}{i,t}$, which is the logged total withdrawal volume from protocol $i$ at time $t$. To control for broader market sentiment, we incorporate $\text{FGI}_{t}$ as a proxy for investor sentiment in the cryptocurrency market and $\text{VIX}_{t}$ for sentiment in the traditional financial market. The sentiment variables are used after logarithmic transformation.
To capture heterogeneity across blockchain layers, protocol versions, and lending platforms, we introduce three dummy variables. To indicate the blockchain layer, we use $\mathrm{Lay}_{i}\in\{0=\text{L1},1=\text{L2}\}$; for the protocol version, we use $\mathrm{Ver}_{i}\in\{0=\text{v2},1=\text{v}3\}$; and for the individual protocol, we use $\mathrm{Platf}_{i}\in\{0=\text{Compound},1=\text{Aave}\}$. In cross-version and cross-chain analysis, we include the dummy variable $\text{Ver}_{i}$ and $\text{Lay}_{i}$ in the regression models—referred to as Model 1 (regression on TVL) and Model 2 (regression on TR). For the analysis of cross-protocol and cross-version effects on performance, we incorporate the dummy variables $\text{Ver}_{i}$ and $\text{Platf}_{i}$, corresponding to Model 3 (regression on TVL) and Model 4 (regression on TR). A comprehensive regression model including all variables is presented in Appendix B.
@@ V Results and Findings
Liq 0.01-0.01 ∗∗∗ $0.00$ $0.00$ $0.00$
Liq:Lay $0.05$ ∗∗∗ $0.09$ ∗∗∗
Liq:Ver $0.03$ ∗∗∗ $0.02$  $0.05$ ∗∗∗ $0.06$ ∗∗∗
Liq:Platf 0.01-0.01  $0.00$
DRM $0.01$ ∗∗∗ $0.01$ ∗∗∗ $0.01$ ∗∗∗ $0.01$ ∗∗∗
VolETH 1.90-1.90 ∗∗∗ 2.23-2.23 ∗∗ 1.94-1.94 ∗∗∗ 2.25-2.25 ∗∗
VolGasPri 0.15-0.15 ∗∗∗ 0.19-0.19 ∗∗∗ 0.15-0.15 ∗∗∗ 0.19-0.19 ∗∗∗
VolUSDCbor $0.02$ ∗∗∗ $0.05$ ∗∗∗ $0.02$ ∗∗∗ $0.05$ ∗∗∗
Wd $0.40$ ∗∗∗ $0.49$ ∗∗∗ $0.40$ ∗∗∗ $0.49$ ∗∗∗
FGI $0.08$ ∗∗∗ $0.46$ ∗∗∗ $0.09$ ∗∗ $0.48$ ∗∗∗
VIX 0.27-0.27 ∗∗∗ 0.76-0.76 ∗∗∗ 0.29-0.29 ∗∗∗ 0.79-0.79 ∗∗∗
@@
V-A Cross-version and cross-chain analysis of liquidations
@@
It should be noted that the coefficient $\mathrm{Liq\mathord{:}Lay}$ is significantly positive at $0.05$, indicating that liquidation has a heterogeneous impact on TVL across different layers under the same version, specifically in v3. This suggests that a $1\%$ increase in liquidation on Arbitrum is associated with a $0.05\%$ greater increase in TVL compared to Ethereum, highlighting a differentiated response to liquidation activity between the two layers. This is a new finding that has not been reported before.
@@ VI Conclusion
We have investigated the effectiveness of automated risk management strategies during liquidation events and their impact on the performance and stability (measured by TVL and TR) of decentralised lending protocols. For this purpose, we selected the top lending protocols, Aave and Compound, including their old (v2) and upgraded (v3) versions deployed on L1 and L2 blockchains, to assess the effectiveness of prevalent risk management mechanisms. The results suggest that liquidation events in v3 protocols lead to increased TVL and total revenue, with a more pronounced effect on the Arbitrum (L2) blockchain compared to the Ethereum (L1) blockchain. In contrast, for the v2 protocols, liquidations have an insignificant impact on TVL and total revenue. A combination of protocol-specific and market-driven factors is likely to influence the behavioral divergence between v2 and v3. Our research has important implications for risk managers, investors, and policy makers.
This research could be further extended by investigating the wallet-level behavior of lenders, borrowers, and liquidators during and after liquidation events to assess strategic user actions. Furthermore, an event study could be used to explore the role of specific stablecoins and volatile assets in causing systemic risk or spillover effects, after cascading liquidations or oracle failure. Another avenue of investigation is to explore the total value redeemable (TVR) as a performance measure alternative to TVL, as it has been proposed as a more accurate quantifier of the value stored within a protocol [11]. In the longer term, our aim is to explore the effects of different risk mechanisms using agent-based simulation modelling [27].
@@ Acknowledgements
J. Cartlidge is supported by UK Research and Innovation (UKRI) Engineering and Physical Sciences Research Council (EPSRC) Grant Number EP/Y028392/1: AI for Collective Intelligence (AI4CI). E. Iftikhar’s research is supported by PNRR (National Recovery and Resilience Plan) scholarship ex D.M. 351/2022 under National Call for PhD in Blockchain and Distributed Ledger Technology offered by University of Camerino (Administrative University) and Universita Cattolica del Sacro Cuore, Milan (Host University). This research was completed at the University of Bristol while E. Iftikhar was a visiting PhD scholar under the supervision of W. Wei and J. Cartlidge.
@@ Appendix A Multicollinearity Test
Variable Liq DRM VolETH VolGasPri VolUSDCbor FGI VIX Wd
VIF 1.25 1.13 1.28 1.61 1.61 2.11 1.84 1.36
@@ Appendix B Comprehensive Model
Here we have used all variable mentioned in our regression model (see Section IV) to account for cross-version, cross-chain and cross-protocol affect on protocol performance (proxied by TVL and TR). Our findings in Table B.2 are consistent with the conclusions we made in Section V. The results show that there is significant heterogeneity between layers and versions, which is consistent with the results of Models 1–4.
@@
Liq:Lay $0.05$ ∗∗∗ $0.09$ ∗∗∗
Liq:Ver $0.03$ ∗∗∗ $0.02$
Liq:Platf 0.01-0.01  $0.00$
DRM $0.01$ ∗∗∗ $0.01$ ∗∗∗
VolETH 1.93-1.93 ∗∗ 2.23-2.23 ∗∗
VolGasPri 0.15-0.15 ∗∗∗ 0.19-0.19 ∗∗∗
VolUSDCbor $0.02$ ∗∗∗ $0.05$ ∗∗∗
Wd $0.40$ ∗∗∗ $0.49$ ∗∗∗
FGI $0.08$ ∗∗ $0.46$ ∗∗∗
VIX 0.27-0.27 ∗∗∗ 0.76-0.76 ∗∗∗